OpenStudy (anonymous):

how do I solve x dy/dx+2y= x^-3

OpenStudy (anonymous):

in standard form I get dy/d\[\frac{ dy }{ dx}+ \frac{ 2y }{ x } = x ^{-4}\]

OpenStudy (anonymous):

Nevermind I got it :D the IF is x^2 then \[x^{2} \frac{ dy }{ dx} +\frac{ 2x^2 }{y}=\frac{ 1 }{ x^2}\] which is then the derivative of \[x^{2}y\] take the integral of both sides and we get \[\int\limits \frac{ 1 }{ x^2 }dx =\frac{ -1 }{ x} +c\] solve for y \[x^2y= \frac{ -1 }{ x} +c\] \[y= \frac{ c }{ x^2}-\frac{ 1 }{ x^3 }\]